OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? -6 + f = 13

OpenStudy (anonymous):

help me work it out and solve it for me please

OpenStudy (anonymous):

Well what can you add to -6 to get rid of it? (make it 0)

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

-6 + -6 = -12

OpenStudy (anonymous):

that's not 0

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

what plus -6 equals 0?

OpenStudy (anonymous):

+6

OpenStudy (anonymous):

indeed.

OpenStudy (anonymous):

So if you add 6 to both sides of the equation what will you have?

OpenStudy (anonymous):

-6 + f + 6 =?

OpenStudy (anonymous):

19i think

OpenStudy (anonymous):

That's what you have on the right side true, but what do you have on the left side of the equation?

OpenStudy (anonymous):

i dont know i really dont

OpenStudy (anonymous):

1

OpenStudy (anonymous):

no. \[-6 + f + 6 \]\[= 6-6 + f \]\[= 0 +f\]\[=f\]

OpenStudy (anonymous):

So when you add 6 to both sides of \(-6+f = 13\), you have \[6 -6 +f = 6+13\] \[f = 19\]

OpenStudy (anonymous):

oh ok